# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v3

@@ Abstract
We establish convergence rates for a fully discrete, multi-level, linear collocation method solving parametric elliptic PDEs on bounded polygonal domains with log-normal inputs. The method uses a finite set of function evaluations in the spatial-parametric domain. Compared with the best-known fully discrete collocation rates, these rates are significantly improved and, up to logarithmic factors, match the rates of best n-term approximations. The results follow from applying general multi-level linear sampling recovery theory in abstract Bochner spaces – via extended least-squares – to infinite-dimensional holomorphic functions. The abstract multi-level recovery in Bochner spaces guarantees yield the improved rates when specialized to the parametric PDE setting.
@@ 1 Introduction and main results
${\rm div}(a({\boldsymbol{y}})\nabla u({\boldsymbol{y}}))\ =\ f\quad\text{in}\quad D,\quad u({\boldsymbol{y}})|_{\partial D}\ =\ 0.$ (1.2)
In the present paper, we consider the case when the diffusion coefficient $a$ is of the lognormal form
$a({\boldsymbol{y}})=\exp(b({\boldsymbol{y}})),\quad{\text{with }}\ b({\boldsymbol{y}})=\sum_{j=1}^{\infty}y_{j}\psi_{j},$ (1.3)
and $y_{j}$ are i.i.d. standard Gaussian random variables, where $\psi_{j}\in L_{\infty}(D)$.
For $r\in{\mathbb{N}}_{0}$ and $\varkappa\in{\mathbb{R}}$, the Kondrat’ev spaces ${\mathcal{K}}^{r}_{\varkappa}(D)$ and ${\mathcal{W}}^{r}_{\infty}(D)$ are defined as the weighted normed spaces of functions on $D$ equipped with norms
$\|u\|_{{\mathcal{K}}^{r}_{\varkappa}(D)}:=\ \sum_{|{\boldsymbol{\alpha}}|\leq r}\|\tau_{D}^{|{\boldsymbol{\alpha}}|-\varkappa}D^{\boldsymbol{\alpha}}u\|_{L^{2}(D)}\qquad\text{and}\qquad\|u\|_{{\mathcal{W}}^{r}_{\infty}(D)}:=\ \sum_{|{\boldsymbol{\alpha}}|\leq r}\|\tau_{D}^{|{\boldsymbol{\alpha}}|}D^{\boldsymbol{\alpha}}u\|_{L^{\infty}(D)},$
@@
By [26, Theorem 3.29], for $r\geq 2$, $f\in{\mathcal{K}}^{r-2}_{\varkappa-1}(D)$ and $a\in{\mathcal{W}}^{r-1}_{\infty}(D)$, the weak solution to (1.1) belongs to ${\mathcal{K}}_{\varkappa+1}^{r}(D)$ provided that with
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{L^{\infty}},\ \ \text{and}\ \ b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\},$
@@
Then for all $n\geq 2$ there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\mathbb{R}}^{\infty}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\begin{split}\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq 1/p_{2}-3/2,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ 1/p_{2}-3/2<\alpha<1/p_{2}-1,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/p_{2}-1,\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/p_{2}-1,\end{cases}\end{split}$ (1.5)
for an arbitrarily small number $\varepsilon>0$, where the constants $C$ and $C_{\varepsilon}$ are independent of $u$ and $n$.
@@
Notice that $S_{n}$ is a fully discrete multi-level collocation approximation method. At each level in $S_{n}$, the spatial component is based on finite element Lagrange interpolation associated with triangulations of the spatial domain $D$, while the parametric component is based on Hermite-Lagrange GPC interpolation in the case $\alpha\leq 1/p_{2}-3/2$, and on extended least squares sampling algorithms in the cases $\alpha>1/p_{2}-3/2$.
@@ 2 Sampling recovery in Bochner spaces
In this section, we investigate multi-level linear sampling recovery of functions in abstract Bochner spaces $L_{2}(U,X;\mu)$ for a Hilbert space $X$ and a probability measure space $(U,\Sigma,\mu)$, based on some weighted $\ell_{2}$-summability of their coefficients of an orthonormal expansion, and some approximation properties in the space $X$. We develop the methods used in [7, 21, 26] to construct fully discrete (multi-level) sampling algorithms by using the approximation properties in the space $X$ combined with extended least squares sampling algorithms at each level, and prove convergence rates of approximation by them. The received results are then applied to $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$, the Bochner space associated with a Hilbert space $X$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$.
@@ 2.1 Function spaces
Let $(U,{\mathcal{A}},\mu)$ be a probability measure space with $\mathcal{A}$} being countably generated and let $X$ be a complex separable Hilbert space. Denote by $L_{2}(U,X;\mu)$ the Bochner space of strongly $\mu$-measurable mappings $v$ from $U$ to $X$, equipped with the norm
@@
Notice that because $\mathcal{A}$} is countably generated, $L_{2}(U,{\mathbb{C}};\mu)$ is separable by [19, Proposition 3.4.5]. Hence $L_{2}(U,X;\mu)$ is a separable complex Hilbert space and, moreover,
@@
Moreover, for every $v\in L_{2}(U,X;\mu)$ represented by the series (2.2), Parseval’s identity holds
@@
Let ${\boldsymbol{\sigma}}=\left(\sigma_{s}\right)_{s\in{\mathbb{N}}}$ be a non-decreasing sequence of positive numbers strictly larger than $1$ such that ${\boldsymbol{\sigma}}^{-1}:=\left(\sigma_{s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{2}({\mathbb{N}})$. For given $U$ and $\mu$, denote by $H_{X,{\boldsymbol{\sigma}}}$ the linear subspace of all functions $v\in L_{2}(U,X;\mu)$ such that the norm
@@
In particular, the space $H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ is the linear subspace in $L_{2}(U,{\mathbb{C}};\mu)$ equipped with its own inner product
@@
having the eigenfunctions $\left(\varphi_{s}\right)_{s\in{\mathbb{N}}}$ and the eigenvalues $\left(\sigma_{1,s}^{-1}\right)_{s\in{\mathbb{N}}}$. Moreover, $K({\boldsymbol{x}},{\boldsymbol{y}})$ satisfies the finite trace assumption
@@
Let $(V_{k})_{k\in{\mathbb{N}}_{0}}$ be a given sequence of subspaces $V_{k}\subset{\color[rgb]{0,0,0}{L_{2}(U,X;\mu)}}$ of dimension $2^{k}$, and $(P_{k})_{k\in{\mathbb{N}}_{0}}$ a given sequence of uniformly bounded linear projectors $P_{k}$ from $\color[rgb]{0,0,0}{L_{2}(U,X;\mu)}$} onto $V_{k}$. For $i=1,2$, let $0<q_{i}\leq 2$ and ${\boldsymbol{\sigma}}_{i}:=(\sigma_{i;s})_{s\in{\mathbb{N}}}$ be given non-decreasing sequences of numbers strictly larger than $1$, such that ${\boldsymbol{\sigma}}_{i}^{-1}:=\left(\sigma_{i,s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{q_{i}}({\mathbb{N}})$. For a given number $\alpha>0$, denote by $H_{X}^{\alpha}$ the linear subspace in $L_{2}(U,X;\mu)$ of all $v$ such that the norm
@@
$\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha}}:=\ \sup_{k\in{\mathbb{N}}_{0}}2^{\alpha k}\left\|{v-P_{k}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}.$ (2.4)
@@
$\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha,\tau}}:=\ \left(\sum_{k\in{\mathbb{N}}_{0}}\left(2^{\alpha k}k^{-\tau}\left\|{v-P_{k}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}\right)^{2}\right)^{1/2}<\infty.$ (2.6)
@@
2.2 Extended least squares sampling algorithms in Bochner spaces
We study the approximate recovery of functions in the space $H_{X,{\boldsymbol{\sigma}}}$ from a finite set of their samples. To ensure a coherent formulation of the problem, we begin with a preliminary observation.
From the separability of $X$ it follows that there exists a set $U_{0}\subset U$ satisfying $\mu(U\setminus U_{0})=0$ such that
and
This means that the pointwise evaluations $f({\boldsymbol{y}})$ are well-defined for every $\boldsymbol{y}$} belonging the full measure subset $U_{0}$ of $U$.
Let $\left(\varphi_{k}\right)_{k\in{\mathbb{N}}}$ be an orthonormal basis of $X$. If $v\in H_{X,{\boldsymbol{\sigma}}}$, then $\langle v,\psi_{k}\rangle_{X}\in H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ for every $k$. Consequently, the pointwise evaluations $\langle v({\boldsymbol{y}}),\psi_{k}\rangle_{X}$ are well-defined for every ${\boldsymbol{y}}\in U_{0}$. This implies that the pointwise evaluations $v({\boldsymbol{y}})$ are also well-defined for every ${\boldsymbol{y}}\in U_{0}$.
Throughout this paper, in the context of sampling recovery, the inclusion $v\in H_{X,{\boldsymbol{\sigma}}}$ means that $v$ is a representative of an element from $L_{2}(U,X;\mu)$ that the pointwise evaluations $v({\boldsymbol{y}})$ for every ${\boldsymbol{y}}\in U_{0}$ are well-defined with $U_{0}$ fixed.
Let us formulate the problem of sampling recovery for $v\in H_{X,{\boldsymbol{\sigma}}}$ as follows: Given sample points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{n}\in U_{0}$ and functions $h_{1},\ldots,h_{n}\in L_{2}(U,{\mathbb{C}};\mu)$, we consider the approximate recovery of $v$ from its values $v({\boldsymbol{y}}_{1}),\ldots,v({\boldsymbol{y}}_{n})$ by the linear sampling algorithm $S_{n}^{X}$ on $U$ defined as
Let $F$ be a subset of $H_{X,{\boldsymbol{\sigma}}}$. Let ${\mathcal{S}}_{n}^{X}$ be the family of all linear sampling algorithms $S_{k}^{X}$ in $L_{2}(U,X;\mu)$ of the form (2.7) with $k\leq n$. To study the optimality of linear sampling algorithms from ${\mathcal{S}}_{n}^{X}$ for the set $F$ and their convergence rates we use the (linear) sampling $n$-width
The following result has been proven in [7, Theorem 2.1] .
Lemma 2.1.
Given arbitrary sample points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{n}\in U$ and functions $h_{1},\ldots,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, for the sampling algorithm $S_{n}^{X}$ in $L_{2}(U,X;\mu)$ defined by (2.7), we have
In the next step, we extend this result to the space $H_{X}^{\alpha,\tau}$. Let $A^{X}$ be a linear operator in $L_{2}(U,X;\mu)$ defined for $v\in L_{2}(U,X;\mu)$ by
@@ 2.2 Extended least squares sampling algorithms in Bochner spaces
Conversely, a linear operator $A^{\mathbb{C}}$ in $L_{2}(U,{\mathbb{C}};\mu)$ of the form (2.9) uniquely defines the $A^{X}$ linear operator of the form (2.8) in $L_{2}(U,X;\mu)$, i.e., this is an one-onto-one correspondence between the linear operators in $L_{2}(U,X;\mu)$ and $L_{2}(U,{\mathbb{C}};\mu)$.

### v4

@@ Abstract
We establish convergence rates for a fully discrete, multi-level, linear collocation method solving parametric elliptic PDEs on bounded polygonal domains with log-normal random inputs. The method uses a finite set of function evaluations in the spatial-parametric domain. Compared with the best-known fully discrete collocation rates, these rates are significantly improved and, up to logarithmic factors, match the rates of best $n$-term approximations. The results follow from applying general multi-level linear sampling recovery theory in abstract Bochner spaces – via extended least-squares – to infinite-dimensional holomorphic functions. The abstract multi-level recovery in Bochner spaces guarantees yield the improved rates when specialized to the parametric PDE setting.
@@ 1 Introduction and main results
${\rm div}(a({\boldsymbol{y}})\nabla u({\boldsymbol{y}}))\ =\ f\quad\text{in}\quad D,\quad u({\boldsymbol{y}})|_{\partial D}\ =\ 0,$ (1.2)
where $a({\boldsymbol{y}})$ is of the log-normal form
$a({\boldsymbol{y}})=\exp(b({\boldsymbol{y}}))\ \ {\text{with }}\ b({\boldsymbol{y}})=\sum_{j=1}^{\infty}y_{j}\psi_{j},$ (1.3)
$y_{j}$ are i.i.d. standard Gaussian random variables and $\psi_{j}\in L_{\infty}(D)$.
[…]
@@
[…]
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\ \ \text{and}\ \ b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\},$
@@
For arbitrary positive number $\varepsilon$, there exist positive constants $C$ and $C_{\varepsilon}$ such that for all $n\geq 2$, there exist points $({\boldsymbol{x}}_{1},{\boldsymbol{y}}_{1}),...,({\boldsymbol{x}}_{n},{\boldsymbol{y}}_{n})\in D\times{\mathbb{R}}^{\infty}$ and functions $\varphi_{1},...,\varphi_{n}\in V$ and $h_{1},...,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{R}};\gamma)$ such that for the linear sampling algorithm $S_{n}$ on the spatial-parametric domain $D\times{\mathbb{R}}^{\infty}$ defined by
@@
$\begin{split}\big\|u-S_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}Cn^{-\alpha}\quad&{\rm if}\ \alpha\leq 1/p_{2}-3/2,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ 1/p_{2}-3/2<\alpha<1/p_{2}-1,\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/p_{2}-1,\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/p_{2}-1.\end{cases}\end{split}$ (1.5)
@@
It is worth emphasizing that $S_{n}$ is a fully discrete multi-level collocation approximation method. At each level in $S_{n}$, the spatial component is based on finite element Lagrange interpolation associated with triangulations of the spatial domain $D$, while the parametric component is based on Hermite-Lagrange GPC interpolation in the case $\alpha\leq 1/p_{2}-3/2$, and on extended least squares sampling algorithms in the cases $\alpha>1/p_{2}-3/2$.
@@ 2 Sampling recovery in Bochner spaces
In this section, we investigate multi-level linear sampling recovery of functions in abstract Bochner spaces $L_{2}(U,X;\mu)$ for a Hilbert space $X$ and a probability measure $\mu$ , based on some weighted $\ell_{2}$-summability of their coefficients of an orthonormal expansion, and some approximation properties in the space $X$. We develop the methods used in [7, 20, 25] to construct fully discrete (multi-level) sampling algorithms by using the approximation properties in the space $X$ combined with extended least squares sampling algorithms at each level, and prove convergence rates of approximation by them. The received results are then applied to $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$, the Bochner space associated with a Hilbert space $X$ and the infinite standard Gaussian measure $\gamma$ on $\mathbb{R}$$^{\infty}$.
@@ 2.1 Function spaces
Let $(U,{\mathcal{A}},\mu)$ be a probability measure space with $U$ being a separable topological space and let $X$ be either a separable complex Hilbert space or $\mathbb{C}$}. Denote by $L_{2}(U,X;\mu)$ the Bochner space of strongly $\mu$-measurable mappings $v$ from $U$ to $X$, equipped with the norm
@@
The space $L_{2}(U,{\mathbb{C}};\mu)$ consists of all square $\mu$-integrable complex functions on $U$. Since $U$ separable, the spaces $L_{2}(U,{\mathbb{C}};\mu)$ and $L_{2}(U,X;\mu)$ are also separable complex Hilbert spaces, and moreover,
@@
Moreover, for every $v\in L_{2}(U,X;\mu)$ represented by the series (2.2), Parseval’s identity holds:
@@
Let ${\boldsymbol{\sigma}}=\left(\sigma_{s}\right)_{s\in{\mathbb{N}}}$ be a non-decreasing sequence of positive numbers strictly larger than $1$ such that ${\boldsymbol{\sigma}}^{-1}:=\left(\sigma_{s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{2}({\mathbb{N}})$. For given $U$ and $\mu$, denote by $H_{X,{\boldsymbol{\sigma}}}$ the subspace of all functions $v\in L_{2}(U,X;\mu)$ such that the norm
@@
In particular, the space $H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ is the subspace in $L_{2}(U,{\mathbb{C}};\mu)$ equipped with its own inner product
@@
[…]
@@
From the separability of $H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ it follows that there exists a set $U_{0}\subset U$ satisfying $\mu(U\setminus U_{0})=0$ such that
and
This means that the pointwise evaluations $f({\boldsymbol{y}})$ are well-defined for every $\boldsymbol{y}$} belonging the full measure subset $U_{0}$ of $U$ (cf. [24]). Let $\left(\psi_{k}\right)_{k\in{\mathbb{N}}}$ be an orthonormal basis of $X$. If $v\in H_{X,{\boldsymbol{\sigma}}}$, then $\langle v,\psi_{k}\rangle_{X}\in H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ for every $k$. Consequently, the pointwise evaluations $\langle v({\boldsymbol{y}}),\psi_{k}\rangle_{X}$ are well-defined for every ${\boldsymbol{y}}\in U_{0}$. This implies that the pointwise evaluations $v({\boldsymbol{y}})$ are also well-defined for every ${\boldsymbol{y}}\in U_{0}$.
Let $X\not$$\mathbb{C}$} and $(V_{k})_{k\in{\mathbb{N}}_{0}}$ be a given sequence of subspaces $V_{k}\subset X$ of dimension $2^{k}$, and $(P_{k}^{X})_{k\in{\mathbb{N}}_{0}}$ a given sequence of uniformly bounded linear projectors $P_{k}^{X}$ from $X$ onto $V_{k}$. We extend the operators $P_{k}^{X}$ as uniformly bounded linear operators in $L_{2}(U,X;\mu)$ (which with an abuse is denoted again by $P_{k}^{X}$) such that
$(P_{k}^{X}v)({\boldsymbol{y}}):=P_{k}^{X}(v({\boldsymbol{y}})),\ \ {\boldsymbol{y}}\in U_{0}.$ (2.3)
For $i=1,2$, let $0<q_{i}\leq 2$ and ${\boldsymbol{\sigma}}_{i}:=(\sigma_{i;s})_{s\in{\mathbb{N}}}$ be given non-decreasing sequences of numbers strictly larger than $1$, such that ${\boldsymbol{\sigma}}_{i}^{-1}:=\left(\sigma_{i,s}^{-1}\right)_{s\in{\mathbb{N}}}\in\ell_{q_{i}}({\mathbb{N}})$. For a given number $\alpha>0$, denote by $H_{X}^{\alpha}$ the linear subspace in $L_{2}(U,X;\mu)$ of all $v$ such that the norm
@@
$\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha}}:=\ \sup_{k\in{\mathbb{N}}_{0}}2^{\alpha k}\left\|{v-P_{k}^{X}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}.$ (2.5)
@@
$\|v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}^{\alpha,\tau}}:=\ \left(\sum_{k\in{\mathbb{N}}_{0}}\left(2^{\alpha k}k^{-\tau}\left\|{v-P_{k}^{X}v}\right\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}\right)^{2}\right)^{1/2}<\infty.$ (2.7)
@@
Let $A^{X}$ be a linear operator in $L_{2}(U,X;\mu)$ defined for $v\in L_{2}(U,X;\mu)$ by
@@
Conversely, a linear operator $A^{\mathbb{C}}$ in $L_{2}(U,{\mathbb{C}};\mu)$ of the form (2.9) uniquely defines the $A^{X}$ linear operator of the form (2.8) in $L_{2}(U,X;\mu)$, i.e., this is an one-onto-one correspondence between the linear operators in $L_{2}(U,X;\mu)$ and $L_{2}(U,{\mathbb{C}};\mu)$. If $A^{X}$ is linear and bounded, then there holds the equality [7, Lemma 2.1]
$\big\|A^{X}\big\|_{H_{X,{\boldsymbol{\sigma}}}\to L_{2}(U,X;\mu)}=\big\|A^{{\mathbb{C}}}\big\|_{H_{{\mathbb{C}},{\boldsymbol{\sigma}}}\to L_{2}(U,{\mathbb{C}};\mu)}.$ (2.10)
These correspondence and norm equality between bounded linear operators $A^{X}$ and $A^{\mathbb{C}}$ allows us to define the spaces $H_{\mathbb{C}}^{\alpha}$ and $H_{{\mathbb{C}}}^{\alpha,\tau}$ generated from given spaces $H_{X}^{\alpha}$ and $H_{X}^{\alpha,\tau}$, by replacing $P_{k}^{X}$ with $P_{k}^{\mathbb{C}}$ and $H_{X,{\boldsymbol{\sigma}}_{i}}$ with $H_{{\mathbb{C}},{\boldsymbol{\sigma}}_{i}}$, $i=1,2$, in their definitions (2.4)–(2.5) and (2.6)–(2.7), respectively. For simplicity, in what follows, we will use the abbreviations $P_{k}:=P_{k}^{X}$ and $P_{k}:=P_{k}^{\mathbb{C}}$, provided this does not lead to any misunderstanding.
We will need the following lemma.
Lemma 2.1.
We have
$\|v-P_{k}v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}\ =\ \left(\sum_{s\in{\mathbb{N}}}\left(\sigma_{2;s}\|v_{s}-P_{k}v_{s}\|_{X}\right)^{2}\right)^{1/2},\ \ v\in H_{X,{\boldsymbol{\sigma}}_{2}}.$
Proof. By the definition,
$\|v-P_{k}v\|_{H_{X,{\boldsymbol{\sigma}}_{2}}}:=\ \left(\sum_{s\in{\mathbb{N}}}\left(\sigma_{s}\|(v-P_{k}v)_{s}\|_{X}\right)^{2}\right)^{1/2}<\infty.$
Due to the equality $(v-P_{k}v)_{s}=v_{s}-(P_{k}v)_{s}$, to prove the lemma it is sufficient to show $(P_{k}v)_{s}=P_{k}v_{s}$. Since ${\boldsymbol{\sigma}}_{2}^{-1}\in\ell_{2}({\mathbb{N}})$, the series (2.2) converges unconditionally in $L_{2}(U,X;\mu)$ to $v$ for $v\in H_{X,{\boldsymbol{\sigma}}_{2}}$. Indeed, by the Hölder inequality we have for $v\in H_{X,{\boldsymbol{\sigma}}_{2}}$,
$\begin{split}\sum_{s\in{\mathbb{N}}}\|v_{s}\varphi_{s}\|_{L_{2}(U,X;\mu)}\ &=\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\|\varphi_{s}\|_{L_{2}(U,{\mathbb{C}};\mu)}\ =\ \sum_{s\in{\mathbb{N}}}\|v_{s}\|_{X}\\ &\leq\bigg(\sum_{s\in{\mathbb{N}}}(\sigma_{s}\|v_{s}\|_{X})^{2}\bigg)^{1/2}\bigg(\sum_{s\in{\mathbb{N}}}\sigma_{s}^{-2}\bigg)^{1/2}\ <\ \infty.\end{split}$
This yields that the series (2.2) absolutely and hence unconditionally converges in $L_{2}(U,{\mathbb{C}};\mu)$ to $v$, since in a Banach space the absolute convergence implies the unconditional convergence. From this unconditional convergence and the uniform boundedness of the linear projectors $(P_{k})_{k\in{\mathbb{N}}_{0}}$ we derive
$P_{k}v\ =\ \sum_{s\in{\mathbb{N}}}P_{k}v_{s}\varphi_{s},$
and therefore, $(P_{k}v)_{s}=P_{k}v_{s}$.
2.2 Extended least squares sampling algorithms in Bochner spaces
We study the approximate recovery of functions in the space $H_{X,{\boldsymbol{\sigma}}}$ or $H_{X}^{\alpha,\tau}$ from a finite set of their samples. Throughout this paper, in the context of sampling recovery, the inclusion $v\in H_{X,{\boldsymbol{\sigma}}}$ or $v\in H_{X}^{\alpha,\tau}$ means that $v$ is a representative of an element from $L_{2}(U,X;\mu)$ that the pointwise evaluations $v({\boldsymbol{y}})$ for every ${\boldsymbol{y}}\in U_{0}$ are well-defined with $U_{0}$ fixed.
Let us formulate the problem of sampling recovery for $v\in H_{X,{\boldsymbol{\sigma}}}$ or $H_{X}^{\alpha,\tau}$ as follows: Given sample points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{n}\in U_{0}$ and functions $h_{1},\ldots,h_{n}\in L_{2}(U,{\mathbb{C}};\mu)$, we consider the approximate recovery of $v$ from its values $v({\boldsymbol{y}}_{1}),\ldots,v({\boldsymbol{y}}_{n})$ by the linear sampling algorithm $S_{n}^{X}$ on $U$ defined as
Let $F$ be a subset of $H_{X,{\boldsymbol{\sigma}}}$ or $H_{X}^{\alpha,\tau}$. Let ${\mathcal{S}}_{n}^{X}$ be the family of all linear sampling algorithms $S_{k}^{X}$ in $L_{2}(U,X;\mu)$ of the form (2.11) with $k\leq n$. To study the optimality of linear sampling algorithms from ${\mathcal{S}}_{n}^{X}$ for the set $F$ and their convergence rates we use the (linear) sampling $n$-width
From (2.10) it follows that [7, Theorem 2.1]
Lemma 2.2.
Given arbitrary sample points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{n}\in{\color[rgb]{0,0,0}{U_{0}}}$ and functions $h_{1},\ldots,h_{n}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\mu)$, for the sampling algorithm $S_{n}^{X}$ in $L_{2}(U,X;\mu)$ defined by (2.11), we have